# A Compute and Communication Runtime Model for Loihi 2

### v1

@@ VII Discussion and future work
There are several limitations that highlight opportunities for future research. This work focuses on a single linear layer and a QUBO solver; a larger variety of kernels and the composition of kernels will provide a more comprehensive assessment of the predictive power of this max-affine model. While exhibiting a tight correspondence with measured runtime, our model also exhibits deviations. These modeling errors can be accommodated for by engineering built-in runtime tolerances in latency-critical applications, and the errors could also be reduced with a more complex runtime model that takes into account finer-grained and dynamic runtime effects beyond those considered in this work. This work directly applies to fall-through mode, where generally one kernel is executing on Loihi 2 at a time and each kernel has the full NoC bandwidth. However, in pipeline mode, where several kernels can run in parallel, further congestion in the NoC can occur. (See [1] for fallthrough and pipeline mode.) Researchers can use the TrafficStats tool to quantify multi-kernel NoC congestion.

### v2

@@ VII Discussion and future work
There are several limitations that highlight opportunities for future research. This work focuses on a single linear layer and a QUBO solver; a larger variety of kernels and the composition of kernels will provide a more comprehensive assessment of the predictive power of this max-affine model. While exhibiting a tight correspondence with measured runtime, our model also exhibits deviations. These modeling errors can be accommodated for by engineering built-in runtime tolerances in latency-critical applications, and the errors could also be reduced with a more complex runtime model that takes into account finer-grained and dynamic runtime effects beyond those considered in this work. This work directly applies to fall-through mode, where generally one kernel is executing on Loihi 2 at a time and each kernel has the full NoC bandwidth. However, in pipeline mode, where several kernels can run in parallel, further congestion in the NoC can occur. (See [1] for fall-through and pipeline mode.) Researchers can use the TrafficStats tool to quantify multi-kernel NoC congestion.